OpenStudy (anonymous):

Simplify the expression: \[\sqrt[5]{96}+\sqrt[5]{3}\]

OpenStudy (anonymous):

i get an answer but i can't get it to come out right...

OpenStudy (anonymous):

for solving this , 96=2^5*3 and hence simplifying u get answer as, \[3\times3 ^{1/5}\]

OpenStudy (anonymous):

why 3 times 3^1/5? why not 3 + 3^1/5 as in the original?

OpenStudy (anonymous):

simplifying the first expression we hav, \[(2*2*2*2*2*3)^{1/5} + 3^{1/5}\] now from the first bracket,u take ou 2 and hence u get \[2\times3^{1/5}+3^{1/5}= 3^{1/5}(2+1)=3\times3^{1/5}\]

OpenStudy (anonymous):

and the 1 comes from where? in the (2+1)?

OpenStudy (anonymous):

oh i think i see it. the 3^1/5 is factored... right?

OpenStudy (anonymous):

yea its is \[2\times3^{1/5}+1\times3^{1/5}\]

OpenStudy (anonymous):

krish, i thought you were leaving. lol

OpenStudy (anonymous):

\[hey\ basket \ math\ , \ do \ u\ know\ latex\ ?\]

OpenStudy (anonymous):

not sure that i do. is that a joke? or a person?

OpenStudy (anonymous):

\[I want\ some\ help\]

OpenStudy (anonymous):

with what?